### v1

# Robust High-Fidelity Quantum Entanglement Distribution over Large-Scale Metropolitan Fiber Networks with Co-propagating Classical Signals

Abstract
The Quantum Internet, a network of quantum-enabled infrastructure, represents the next frontier in telecommunications, promising unprecedented capabilities that cannot be attained by classical counterparts. A crucial step in realizing such large-scale quantum networks is the integration of entanglement distribution within existing telecommunication infrastructure. Here, we demonstrate a real-world scalable quantum networking testbed deployed within Deutsche Telekom’s metropolitan fibers in Berlin. Using commercially available quantum devices and standard add-drop multiplexing hardware, we distribute polarization-entangled photon pairs over dynamically selectable fiber paths ranging from 10 m to 82 km. Quantum signals, transmitted at 1324 nm (O-band), coexist with conventional bidirectional C-band traffic without dedicated fibers or infrastructure changes. Active stabilization of the polarization enables robust long-term performance, achieving entanglement fidelities between 85–99% and Clauser-Horne-Shimony-Holt parameter $S$-values between 2.36–2.74 during continuous multiday operation. By achieving high-fidelity entanglement distribution under real urban conditions with less than 1% downtime, we confirm the feasibility of hybrid quantum-classical networks under real-world conditions at metropolitan scale. These results establish deployment benchmarks and provide a practical roadmap for telecom operators to integrate quantum capabilities.
I Introduction
The rapidly advancing field of quantum technology promises revolutionary applications ranging from quantum computing [1, 2, 3] to enhanced metrology [4, 5, 6, 7] and secure communications [8, 9, 10, 11, 12, 13], which will impact diverse areas such as financial transactions [14] and voting systems [15]. At the core of many of these applications lies entanglement, a distinctive quantum correlation that can be spatially separated, providing the foundation for quantum networks often referred to as the Quantum Internet. Entanglement-based networks underpin key future capabilities such as quantum repeaters [16], distributed sensing [7], quantum computing [17], and time synchronization [18].
Photons, ideal carriers for this quantum information due to their weak environmental coupling, will require distribution through existing fiber optics to enable future quantum technologies. However, distributing entangled states over fiber-based telecom infrastructure that already transmits classical data is nontrivial. Key technical hurdles include the development of field-deployable quantum hardware, compatibility with legacy systems, and mitigation of noise, environmental fluctuations, and channel crosstalk. Prior demonstrations of quantum-classical coexistence have mostly been limited to single-channel [19, 20, 21] or controlled laboratory settings [22, 23, 24], with minor emphasis on long-term stability or scalability in real-world or multi-user metropolitan networks. Recently, more realistic approaches are starting to be explored [25, 26, 12].
Figure 1: The BearlinQ layout. Setup schematic for adaptive routing of polarization-entangled photon pairs between local (T-labs) and deployed fiber paths, shown in the lower left with a schematic of DT’s Berlin fiber network nodes from the Winterfeldtstrasse laboratories (WFD). Fibers are used individually or stitched to reach $\sim 15$ km (HHI), $\sim 30$ km (Dotti), $\sim 60$ km (2$\times$ Dotti), or $\sim 82$ km (Dotti & HUB). Polarization-entangled photons are generated at the SRC device optically driven by the LSR. The idler photon ($795,\text{nm}$) is routed to the measurement station, while the signal photon ($1324,\text{nm}$) travels through the APC-INJ/CMP stack, fiber-gapped by the intervening path-length segments, with the path selected by an optical switch. OADMs after the APC-INJ and before the APC-CMP allow co/counter-propagation of C-band channels. After the APC-CMP, signal photons go to the measurement station. A high-pass filter suppresses classical leakage at the SNSPD. Coincidences at given polarizations are logged using a Swabian Instruments TDC, while CHSH measurements are enabled by rotating $\lambda/2$ plates and polarizing beam splitters (PBS).
Quantum network deployments hinge on reliable instrumentation to generate, transmit, and measure entanglement as well as on systems to stabilize the overall infrastructure. A key design choice is the photonic degree of freedom, the most popular examples being time-bins [27], frequency-bins [13], or polarization [28, 29]. Among these, polarization encoding is particularly appealing because of its natural compatibility with other quantum systems such as quantum memories. However, photon polarization is highly sensitive to birefringence fluctuations which degrade entanglement fidelity and protocol reliability, an issue magnified in multi-path topologies common to real-world networks. This makes active polarization compensation essential for stable entanglement distribution.
We demonstrate the integration of stable, high-fidelity polarization entanglement into an existing classical network under the city of Berlin. We call this hybrid quantum-classical network installation BearlinQ, a name reflecting our prototype’s technical robustness while evoking the strength and resilience symbolized by Berlin’s iconic bear. This is attained by distributing polarization-entangled O-band (1324 nm) photons over Deutsche Telekom’s (DT) metropolitan fiber network in the presence of co- and counter-propagating C-band (1550 nm) classical traffic under real-world conditions. Robust quantum performance is enabled by Qunnect’s integrated hardware suite comprised of an entanglement source, pumping lasers, and automatic polarization compensation to dynamically stabilize polarization across deployed fiber links. Integration with existing network infrastructure is made possible by optical add-drop multiplexers (OADMs), dynamic path switching, and adherence to International Telecommunication Union - Telecommunication Standardization Sector (ITU-T) channel standards, all without the need for dedicated fibers. To demonstrate network flexibility, we perform drift and switching experiments on different fiber paths. Long-term stability is tested for multiple days by alternating between local and deployed fiber segments of up to $\sim$82 km. Clauser-Horne-Shimony-Holt (CHSH) inequality [30] and fidelity measurements are periodically performed, confirming strong, continued quantum correlations in all scenarios. With automatic compensation, we maintain high fidelity with long-term quantum stability, demonstrating that reliable and scalable quantum networking is achievable by operators such as DT on existing telecom infrastructure, bringing quantum networking closer to real-world deployment.
The paper is organized into two main sections outlining the network implementation and performance. Section II describes the BearlinQ fiber testbed, covering key hardware elements such as the entanglement source, polarization stabilization, and quantum-classical signal integration. Section III presents experimental results, including stability tests over urban distances, path-switching effects, and key quantum operational metrics for network operators.
II Setup for Urban Entanglement Distribution
II.1 The BearlinQ network
Our quantum-classical hybrid network is centered around an entanglement generation hub connected to deployed fibers. A schematic of the layout is shown in Figure 1. Entanglement is generated by an off-the-shelf commercial device (SRC, see Section II.2) producing entangled photon pairs at 795 nm (NIR, idler) and 1324 nm (telecom O-band, signal), driven by a dedicated laser system (LSR). These rack-mounted (2U) devices enable deployment and comply with telecom collocation standards. The 795 nm photons are routed through local fibers directly to a measurement station. The 1324 nm photons are directed to the same measurement station either along a short local fiber or through deployed fibers that exit the laboratory via a series of patch panels, traverse the BearlinQ network, and loop back to the lab. A 2$\times$2 mechanical optical cross-switch allows dynamic selection of the fiber paths. Classical communication signals required for coordination are transmitted over the C-band and combined with/separated from the quantum channel using OADMs. A central control server orchestrates the experiment and analysis.
To compensate for polarization drifts in the fibers caused by environmental fluctuations or switching events, an automatic polarization compensator (APC) actively maintains the polarization state of the O-band photons. Following entanglement generation, the quantum signal is sent through the device pair, which interweaves classical probing light into the fiber to enable multi-state polarization tracking and applies dynamic corrections to mitigate polarization drifts by time-multiplexing to ensure that any arbitrary polarization will be preserved during long-haul propagation (see Section II.3).
At the measurement station, O-band and NIR photons are analyzed. Detection is performed using superconducting nanowire single-photon detectors (SNSPDs). Arrival times are time-stamped with a time-to-digital converter (TDC) and processed at the server to extract the CHSH $S$-value. The measurement station, consisting of motorized half-wave ($\lambda/2$) plates, fixed polarizing beam splitters, and a bandpass filter to suppress C-band noise, supports automated CHSH testing. Manual controllers fine-tune the local polarization reference and ensure alignment with the SNSPDs. This fine-tuning is required after any setup modification. Run-to-run variations in the measured $S$-values are primarily attributed to uncertainties in this manual calibration procedure (see Supp. Mat. IV.1).
Finally, we characterize our network by performing optical time domain reflectometer (OTDR) measurements on each deployed length of fiber. Explicit characterization is essential as real-world fibers are subject to a range of imperfections including environmental influences, splices, connector interfaces, and inconsistencies in manufacturing or installation which can introduce losses exceeding nominal specifications. This is evident in BearlinQ, where the 30 km and 60 km loops exhibit losses close to those expected, but the 15 km and 82 km loops contain segments with larger attenuation (see Figure 2).
Figure 2: Fiber characterization. OTDR measurements for selected fibers of $\sim$ 15, 30, 60, and 82 km. Trace discontinuities correspond to fiber splices, connectors, and mating sleeves. For reference, the expected attenuation of standard single-mode fiber (SMF-28) at 1324 nm is $\sim 0.32$ dB/km [31], as indicated by the black dashed line.
II.2 Quantum-Classical Integration and Entanglement Source
Important considerations for the design of hybrid networks that need to balance the requirements of both quantum operations and everyday classical use are the choice of relative wavelengths for the classical and quantum signals and the intimately-related selection of an entanglement source. The former is mainly informed by two elements: channel density and noise management. Transmitting quantum and classical signals respectively in the O- and C-bands effectively addresses both issues. Additionally, the selected O-band entanglement source meets necessary reliability benchmarks for stable real-life operation.
The high channel density of the C-band supports a large number of simultaneous users through a mature and standardized dense wavelength division multiplexing (DWDM) infrastructure with narrow channel spacings of 50 GHz. In contrast, the O-band is limited to coarse WDM (CWDM) standards, typically offering only three usable channels with 20 nm spacing, which substantially limits its multiplexing capacity. This disparity naturally assigns classical signals to the C-band, where high data throughput and multi-service access are essential for existing network users, while reserving the O-band for low-noise quantum channels.
Single-photon signals are extremely susceptible to noise in the form of uncorrelated photons that make their way into the measurement devices. The dominant noise source is spontaneous Raman scattering [32], which occurs when light interacts with its propagation medium generating broadband frequency-shifted photons that overlap with the quantum signal and thus cannot be filtered. The accurate selection of the relative wavelengths of the quantum and classical channels effectively mitigates this issue. Raman scattering is an asymmetric process, predominantly producing Stokes-shifted (longer wavelength) photons. Transmitting quantum signals in the O-band and classical data in the C-band is particularly favorable. By placing classical channels at longer wavelengths, the majority of Raman noise falls outside the quantum detection window, significantly avoiding in-band noise pollution. This approach enables the use of high classical launch powers (¿10 dBm), relaxes spectral filtering requirements (reducing insertion loss), and enhances system robustness for real-world deployment of coexisting channels over existing fiber infrastructure.
Having identified the O-band as the ideal range for quantum signals, we employ Qunnect’s SRC as the source of entangled photons. The SRC operates at room-temperature delivering ¿$10^{7}$ photon pairs/s with $\sim 1$ GHz linewidth (high brightness) and intrinsically stable polarization entanglement with large signal-idler cross-correlations ($g^{(2)}_{si}\gtrsim 30$) [33] and a heralding efficiency of $\sim 30\%$ [28]. Its O-band wavelength also leverages the fibers’ 1310 nm low-dispersion window and removes the need for frequency conversion. These features make the SRC an ideal tool for the generation of entanglement in quantum networks run by telecom operators, as it eliminates many constraints of conventional sources such as limited brightness or fidelity, or the need for cryogenics, vacuum apparatus, or controlled laboratory environments.
The device uses diamond-scheme spontaneous four-wave mixing (SFWM) in a warm rubidium 87Rb vapor cell to generate bichromatic polarization-entangled photon pairs [28]. The LSR 780 nm pump and 1367 nm coupling lasers drive the atomic transition from the ground state, $|5S_{1/2}\rangle$, to the doubly excited state $|6S_{1/2}\rangle$. Decays to the $|5P_{1/2}\rangle$ and $|5S_{1/2}\rangle$ states emit photons at 1324 nm and 795 nm, respectively. The phase-matching condition ensures strong temporal and spatial correlations between photon pairs. By matching the polarization of the pump and coupling fields, the Rb spectroscopic structure guarantees the emission of photon pairs in the maximally entangled Bell state |$\ket{\Phi_{+}}=\frac{1}{\sqrt{2}}(\ket{HH}+\ket{VV})$.
The photon pair generation rate scales linearly with pump laser power, as the probability of pair creation is directly proportional to the number of pump photons available for the nonlinear interaction [28]. Increasing the pump power is particularly useful in compensating for losses that accumulate with increasing fiber length, as a higher pair-generation rate helps maintain sufficient photon rates for use-cases. However, while increasing pump power helps overcome larger transmission losses, it also introduces a trade-off in signal-idler correlation strength, as quantified by the cross-correlation function $g^{(2)}_{si}$ [33]. The inverse relationship between pair-rate and $g^{(2)}_{si}$ imposes a practical limitation, as large pump powers ultimately reduce the proportion of usable entangled pairs. To accommodate different fiber lengths while ensuring high fidelity a constant coupling power of $8.4$ mW was used, along with variable pump powers of $\sim 450$ µW and $\sim 900$ µW.
II.3 Active Polarization Stabilization
Environmental factors such as temperature fluctuations [34] and mechanical stress [35] on single-mode, non-polarization maintaining optical fibers induce birefringence changes, causing unwanted rotations of the state of polarization of propagating photons. Protocols relying on polarization require precise control of these states to correctly analyze quantum correlations. Uncompensated variations can effectively randomize the polarization basis, degrading the fidelity and usability of entanglement across the network. To ensure accurate correlation measurements, we actively track and compensate for polarization drifts by maintaining a stable and distributed polarization reference across all nodes. The APC system (injector, APC-INJ, and compensator, APC-CMP) provides the reference and stabilizes the fibers to it. The APC-INJ transmits a known control sequence of approximately the same wavelength as the qubits to the APC-CMP at the opposite end of a fiber link (see Supp. Mat. IV.1). The APC-CMP then analyzes the deviations of a sequence of states from their intended polarizations and applies corrections through a low-loss elasto-electro-optical feedback mechanism [29]. This system uses time-multiplexing to operate over the same fibers as both quantum and classical signals.
Figure 3: Passive fiber operation. Time evolution of $S$-value and quantum fidelity, $F_{Q}$, for a local (blue) and a 60 km deployed (magenta) fiber path, both without active stabilization control. The red dashed line indicates the CHSH threshold at $S=2$, the classical limit for local hidden variable theories. A value $S>2$ is required to certify entanglement and enable quantum applications. A high quantum fidelity (high similarity of the transmitted state to a well-defined reference state) is also necessary for channel suitability, but is not guaranteed by $S>2$. The local fibers within the T-Labs environment exhibit a very stable $S$ and $F_{Q}$ over $\sim 60$ hours. In contrast, the outside path exhibits a sudden drop in both below usable thresholds. Error bars represent $1\sigma$ uncertainties from Poissonian count statistics propagated through CHSH calculations. The integration time for each waveplate configuration was 1 s (2 s) for the local (60 km) path.
Figure 4: CHSH measurements across BearlinQ. Normalized coincidence counts as a function of signal waveplate angle for idler waveplate settings of $0^{\circ},22.5^{\circ},45^{\circ}$, and $67.5^{\circ}$ in four configurations: (a) local fiber loop without the optical switch nor the APC; (b) 60 km fiber loop including a switch and patch panels without the APC; (c) local fiber loop including the APC, the switch, and the patch panel; and (d) 60 km fiber loop including a switch, patch panels and the APC. Error bars represent $1\sigma$ uncertainties from Poissonian count statistics propagated through the CHSH $S$ calculation. The integration time for each waveplate configuration was 1 s (2 s) for the local (60 km) path.
We characterize the fibers to assess drift rate and evaluate the required period of basis re-alignments by performing CHSH and quantum fidelity, $F_{Q}$, measurements over both local and deployed fibers in the absence of active stabilization. The CHSH $S$-value is calculated using standard methods [36], while the upper/lower quantum fidelity bound calculation is detailed in Supp. Mat. IV.2. As shown in Figure 3, the fiber segments flanking the APC pair (common to both the local and deployed paths) and the local T-Labs fiber linking the two APCs exhibit high stability over multiple days. In contrast, while the laboratory environment effectively suppresses drifts, the deployed loop can be subject to instability and unpredictability. In the experimental run presented here, the selected fibers demonstrate resilience to slow drifts but exhibit sudden jumps. This behavior is not consistent across deployments: in other instances, gradual drifts are more prevalent [37]. This variation likely arises from differences in installation practices: specifically, here all fibers are buried, and the infrastructure is modern and actively maintained by DT. Nevertheless, we confirm that active stabilization is required for stable network operation by telecom operators.
Refer to captionRefer to caption
Figure 5: Sustained automated multi-path switching in a metropolitan network. a) Sequence of logic operations showing key timing intervals for CHSH measurement (blue, $t_{\mathrm{CHSH}}$), active polarization control (green, $t_{\mathrm{APC\,comp}}$), path switching (yellow, $T_{\mathrm{path\,switch}}$), and local acquisition (yellow, $t_{\mathrm{local}}$). Vertical dotted lines mark switching events. b) Start and end classical fidelity for all APC routines during the switch from BearlinQ to local fiber. Histograms at right show the distribution of final classical fidelities measured by the APC. c) Local long-term CHSH and quantum fidelity traces, each corresponding to a switch from one of the deployed fibers (see following panels); similar color tones indicate the same measurement run. d), e), f) Long-term CHSH and quantum fidelity over 15 km, 30 km, and 60 km, respectively. Error bars in c–f show $1\sigma$ uncertainties from Poissonian count statistics propagated through the CHSH $S$ calculation. In c–f, the red dashed line marks $S=2$. The integration time for each waveplate configuration in the CHSH and fidelity measurements was 1 s for the local, 15 km, and 30 km paths, and 2 s for the 60 km path.
III Entanglement Distribution Through Berlin
III.1 CHSH over Local & Urban Distances
To characterize our quantum network operation, we measure the CHSH inequality violation and the fidelity of our quantum states relative to the |$\ket{\Phi_{+}}$ Bell state. We establish a baseline for the performance of the entanglement distribution by performing tests both locally and over a 60 km fiber link without APC or path routing. The local CHSH test yields an $S$-value of $S=2.71\pm 0.01$ and a quantum fidelity of $96.84\%\leq F_{Q}\leq 98.31\%$ (Figure 4 a). The same measurement was performed through the switch and over 60 km resulting in a $S$-value of $S=2.60\pm 0.04$ and $95.31\%\leq F_{Q}\leq 96.52\%$ (Figure 4 b). The somewhat worse performance of the 60 km loop is due to the increased SRC pair rate necessary to overcome the fiber losses, as explained in Section II.2.
Flexible path routing is necessary for a truly hybrid network operation that is in accordance with current classical networks run by telecom operators. To enable routing, we introduce automated path-switching in our setup. The APC re-aligns polarization after each path change, ensuring high entanglement fidelity regardless of the route. To confirm that the additional hardware does not negatively affect the results, we repeat the measurement locally with all the devices in place. An $S$-value of $S=2.72\pm 0.01$ is measured, consistent with the initial local results (Figure  4 c). After switching to the 60 km path and running a compensation routine, we measure $S=2.65\pm 0.10$ and quantum fidelity of $96.34\%\leq\text{F}_{Q}\leq 98.38\%$ (Figure  4 d), once again lower due to higher pair rates. These results demonstrate that automated path-switching and polarization compensation function as intended, eliminating the need for manual adjustments and enabling stable quantum correlations over independent fiber paths.
Reference Location Setting Fiber (km) Fidelity (%) Pair Rate (pairs/s) Quantum $\lambda$ (nm) Classical $\lambda$ (nm) Duration This work Berlin, DE Urban $10^{-2}$ to 82 99-85 107- $5\times 10^{2}$ 795, 1324 1560.61 days (ea.) [22] Chicago, US Laboratory 30.2 $>90$ 1290-1310 1547 - [23] Geneva, CH Laboratory 50 $\sim 80$ $\sim 10^{1}$ 1310 1534 hours [24] Singapore, SG Laboratory 50 92.5-97 $10^{4}$ 586, 1310 - - [29] Brooklyn, US Urban 34 99 $5\times 10^{5}$ 795, 1324 - 15 days [19] Stuttgart, DE Urban 35.8 94.5 $\sim 10^{3}$ 780, 1515 - days [20] Chicago, US Urban 47.9 $\sim 99$ $\sim 10^{2}$ 1280-1318 1549-1565 - [21] Shandong, CN Urban 66 N/A N/A 1310 1490, 1538 -
Table 1: Comparison of performance metrics of O-band quantum entanglement distribution or comparable testbeds [^2] To the best of the author’s knowledge these represent the state-of-the-art of O-band or bichromatic photon-pair quantum entanglement distribution. We distinguish two types of experiments: fiber spool-based in-lab setups (incl. between nearby buildings supplemented with spools) (laboratory); real-world deployments using installed fiber (urban). If a metric was unavailable, we used “-”; if not relevant, we used “N/A”.
III.2 Sustained Automated Path Switching
A second benchmark to establish BearlinQ as a mature telecom network is the demonstration of long-term reliable operation. This is shown by continuously operating the system over multiple days with automated switching across available paths. To assess the robustness of the adaptive routing, periodic CHSH and fidelity measurements are performed while alternating the photon route between the local T-lab path and the long-distance fibers every 30 minutes. Each time the fiber path is switched, the APC’s compensation routine is triggered. A time diagram of the routine is shown in Figure 5 a. With a more advanced optical switch and access to additional fiber paths, this approach could scale to route polarization qubits from multiple entanglement sources to various nodes across a metropolitan-scale quantum network.
We run long-term measurements for 15 km, 30 km, 60 km, and 82 km paths over durations ranging from 16 to 65 hours. The 15 km, 30 km, and 60 km measurements are performed in parallel with periodic switching between configurations and active compensation dynamically adapting to varying fiber conditions. Figure 5 b shows the classical APC-calculated start fidelity (pre-compensation) and end fidelity (post-compensation) for each switching event to illustrate the effectiveness of polarization compensation. We see that different fiber paths require distinct polarization transformations, resulting in variable compensation times. This comes at the cost of network downtime, which ranges from tens of milliseconds to a few seconds (average compensation time of 511 ms across all paths of Figure 5). The $95^{\text{th}}$ percentile compensation times for a change from local to 60 km path and vice versa are $9.5$ s and $250$ ms, respectively. Figure  5 c, d, e, and f and Figure 6 show the calculated $S$-values and quantum fidelity bounds for each fiber length. We observe sustained quantum correlations across all configurations, with an average $S=2.66\pm 0.04$ and quantum fidelity bounds between $90\%\leq F_{Q}\leq 99\%$ whenever measuring on the local path. For the 15 km, 30 km, 60 km loops, we measure average CHSH parameters of $S=2.53\pm 0.05$, $S=2.61\pm 0.06$, and $S=2.61\pm 0.09$, respectively. On the 15 km path, the quantum fidelity lower bound exceeded $90\%$ during 75% of the runtime, while on the 30 km and 60 km fibers this threshold was upheld for 94% and 79% of the runtime, respectively. All paths exceed $85\%$ quantum fidelity lower bound $100\%$ of the time, except the 60 km path which does 98.5% of the time. Given the necessity of longer integration times, the 82 km data was taken without path switching. In this longest fiber path, we measure $S=2.58\pm 0.03$ and a fidelity $89.5\%\leq F_{Q}\leq 98.4\%$. All these measurements highlight robust long-term quantum operations over all fiber routes.
Figure 6: Entanglement distribution over 82 km. The results confirm the system’s ability to preserve strong quantum correlations despite high losses ($-42$ dB) and uncontrolled environmental perturbations introduced by the fibers. The integration time for each waveplate configuration in the CHSH and fidelity measurements was 60 s.
III.3 Operational Metrics
Network operators rely on a set of well-defined key performance indicators (KPIs) to assess critical daily services. In conventional telecom networks, these include high uptime (¿99.99%), non-interference between channels, and predictable signal integrity. Despite it being particularly challenging, translating these standards to the quantum domain is necessary to establish commercial hybrid quantum-classical networks. Achieving telecom-grade reliability in the quantum domain requires active stabilization, real-time diagnostics, and extremely low-noise operation, none of which are trivial with entangled photons. By addressing these issues, our system takes a step toward bridging the gap between quantum networks as a research tool and as real-world infrastructure.
BearlinQ meets telecom-inspired KPIs by demonstrating uninterrupted entanglement distribution while routing across metropolitan infrastructure under real-life conditions. Operator-aligned metrics such as low downtime, minimal impact on classical traffic, high rates, and high-fidelity qubit transmission are achieved. In this context, downtime is defined as periods when $F_{Q}$ falls below $85\%$ or the system requires calibration or maintenance. Automatic stabilization ensures near carrier-grade reliability, as demonstrated by $1.5\%$ downtime on the 60 km path while maintaining quantum fidelity $F_{Q}>85\%$ (Figure 5 d). Quantum fidelity was primarily limited by the high entanglement generation rate of the SRC and residual detuning in the APC (see Supp. Mat. IV.2). An additional minor reduction of $-0.3\%$ was attributed to polarization-dependent loss (PDL) totaling $\sim 0.9$ dB, excluding long-haul links, mainly arising from the optical switch and WDM components.
Beyond classically-inspired KPIs, hybrid networks have quantum-specific constraint. As explained in Sec.  II.2, noise in the quantum channel is one of them. To minimize unwanted photons at the quantum wavelengths, our network employs a quantum signal on the 1310 nm CWDM channel mixed using OADMs with classical data on C-band DWDM channels at 1551.32 nm (Ch 27, +6 dBm) and 1560.61 nm (Ch 13, +4 dBm). The results in Figure 7 demonstrate low noise levels in the 60 km path configuration, enabling effective copropagation.
Figure 7: Noise pollution analysis on the 60 km link. Average of correlation peaks for all waveplate configurations in a full CHSH measurement with (magenta) and without (blue) co-propagating classical light. The uncorrelated background noise, primarily from detector dark counts and the SRC’s thermal emission, yields SNRs of 95 (classical off) and 90 (classical on). The classical channel introduces $\sim 30$ additional counts/s on average.
Practical quantum networking also demands useful pair-rates. Instrumental for this result was the high SRC pair-generation rate of $\geq 10^{7}$ pairs/s. Our deployed system consistently delivers coincidence rates ranging from $5\times 10^{2}-10^{7}$ entangled pairs/s after losses, depending on the selected fiber path. Even at attenuation levels of up to $-42$dB (82 km path), our network had $\geq 10^{2}$ pairs/s, showcasing a resilient platform for real-world useful quantum networking. However, in the absence of ultralow-loss fibers or quantum repeaters, the attenuation of the O-band fiber at $\sim$0.32 dB/km limits practical point-to-point deployments to distances of $\lesssim$100 km. Prior O-band demonstrations relied on fiber spools in laboratory settings [23, 24] or short campus loops [25]. Here, we exceed those benchmarks using 82 km of urban fiber despite added insertion losses from the adaptive routing and non-ideal deployed fibers. Table 1 summarizes key performance metrics and highlights how this deployment establishes a new operational benchmark for metropolitan quantum networks.
IV Conclusion
Practical quantum networking requires critical components including bright, deployable entanglement sources producing high-fidelity photon pairs, rapid reference frame stabilization, and low-loss routing. Here, we addressed these challenges and successfully deployed BearlinQ, a quantum testbed in the metropolitan fiber network of Deutsche Telekom in Berlin. Using commercially available devices and telecom-standard optical add-drop multiplexing hardware, we demonstrated uninterrupted high-rate entanglement distribution through dynamically selectable fiber paths ranging from 10 m to 82 km with quantum signals in the telecom 1324 nm O-band seamlessly coexisting with classical bidirectional C-band traffic. Using active polarization stabilization, our system maintained high entanglement fidelity, with lower and upper bounds ranging from 85% to 99% and with S values between 2.36 and 2.74, indicating strong quantum non-locality, with near telecom-grade reliability ($\sim 1$% downtime) over extended durations, all without requiring dedicated fiber or modifications to existing telecom infrastructure. Our deployment also validates the feasibility of quantum routing within existing telecom infrastructure, demonstrating telecom operational reliability requirements over typical metropolitan distances.
By leveraging co-existence of usable quantum photons with standard classical telecom traffic, this work sets operational benchmarks necessary for commercially viable quantum services, eliminating costs and complexity of new fiber installations. These results highlight that quantum networks can be integrated with existing assets and are ready to transition from lab-based demonstrations to reliable deployments. By embracing these innovations, operators can establish themselves as the backbone of quantum-enabled networks, paving the way toward commercially viable quantum services for quantum communications, distributed quantum computing, and quantum sensing applications integrated into existing urban infrastructure.
Supplementary Material
IV.1 Performance Optimizations
CHSH consistency: The calibration procedure begins by inserting perpendicularly aligned polarizers in the SRC to preselect the polarization state |$\ket{H}$ or |$\ket{V}$. Motorized $\lambda/2$-waveplates are then aligned with their slow/fast axes set to the zero angle position, which is then stored as a home reference. Next, manual polarization controllers before the measurement station are adjusted to minimize single-photon counts, thus defining the reflected basis of the PBS’s as the reference basis of measurement. Subsequently, the waveplates are rotated to the diagonal basis where LCR voltages are then fine-tuned to minimize coincidences. These steps could be automated using programmable fiber paddles, or the APC itself, eliminating user-induced variability and enabling more precise and repeatable calibrations than those presented herein.
Fidelity improvements: High-fidelity entanglement distribution relies on spectrally narrowband photons to minimize internal wavelength spread and associated polarization dispersion. However, quantum fidelity is fundamentally constrained when the APC is calibrated at a wavelength different from that of the quantum signal because fiber-induced polarization rotations are both wavelength-dependent and time-varying. Measurements in deployed telecom fibers show that the fiber’s polarization transformation varies significantly with wavelength. This wavelength dependence becomes pronounced over tens of kilometers, rendering a system calibrated at one wavelength inherently misaligned for others. When polarization compensation is performed at 1324 nm, high fidelity is maintained only for photons within a narrow spectral range (¡0.5 nm). As shown in Figure 8, even slight detuning (or a broad photon bandwidth) introduces measurable infidelities. Our data show that spectral mismatch consistently reduces Bell-state fidelity across all fiber configurations. This reduction was the high fiber attenuation limiting APC operation at the target wavelength due to insufficient optical power, leading to a quantum fidelity losses of up to 4%. This constraint can be easily mitigated in future implementations.
Figure 8: APC wavelength impact on quantum fidelity. As probing wavelength deviates from that of qubits, the classically inferred fidelity becomes increasingly unreliable. The discrepancy grows with spectral detuning, underscoring the importance of a near-zero offset to ensure accurate $F_{Q}$. This is further exacerbated by PMD effects in long fibers.
Operational downtime: To minimize network interruptions, we plan to implement continuous monitoring of polarization drift by analyzing a tap on existing co-propagating C-band light with a 15XX nm APC-CMP. This unit would then take advantage of its embedded polarimeter to detect SOP discontinuities, which would be used to trigger the 13XX nm APC. This stabilization protocol eliminates the need to rely solely on routine classical fidelity checks in the time-multiplexed compensation scheme, thereby reducing network downtime.
IV.2 Fidelity Calculations
Despite not performing a full quantum state tomography [38], we can still estimate approximate quantum fidelity bounds based on the assumption that the generated state is close to a maximally entangled bell-state: |$\ket{\phi_{+}}=\frac{1}{\sqrt{2}}(\ket{HH}+\ket{VV})$. The fidelity between the reconstructed density matrix and the target state is given by
$F$ $=(Tr[\ket{\phi_{+}}\bra{\phi_{+}}\rho\ket{\phi_{+}}\bra{\phi_{+}}])^{2}$ (1)
$=\>\frac{\rho_{11}+\rho_{44}+\rho_{41}+\rho_{14}}{2}$ (2)
[…]
$\rho_{11}=\frac{C_{HH}}{N},\;\rho_{44}=\frac{C_{VV}}{N}$ (3)
Coincidence measurements in the diagonal and anti-diagonal basis relate to the off-diagonal elements by
$\frac{C_{DD}+C_{AA}}{N}=\frac{1+\rho_{14}+\rho_{23}+\rho_{32}+\rho_{41}}{2}$ (4)
Substituting this into 2 gives
$F=\frac{C_{HH}+C_{VV}+2C_{DD}+2C_{AA}}{2N}-\frac{1+\rho_{23}+\rho_{32}}{2}$ (5)
We can bound the final term by noting that $|\rho_{23}|=|\rho_{32}|\leq\sqrt{\rho_{22}\rho_{33}}$ such that we obtain
$\begin{array}[]{c}\displaystyle\frac{C_{HH}+C_{VV}+2C_{DD}+2C_{AA}-2\sqrt{C_{HV}C_{VH}}}{2N}-\frac{1}{2}\\[10.00002pt] \leq F\leq\\[5.0pt] \displaystyle\frac{C_{HH}+C_{VV}+2C_{DD}+2C_{AA}+2\sqrt{C_{HV}C_{VH}}}{2N}-\frac{1}{2}\end{array}$ (6)
Similarly, we obtain the second inequality
$\begin{array}[]{c}\displaystyle\frac{C_{DD}+C_{AA}+2C_{HH}+2C_{VV}-2\sqrt{C_{AD}C_{DA}}}{2N}-\frac{1}{2}\\[10.00002pt] \leq F\leq\\[5.0pt] \displaystyle\frac{C_{DD}+C_{AA}+2C_{HH}+2C_{VV}+2\sqrt{C_{AD}C_{DA}}}{2N}-\frac{1}{2}\end{array}$ (7)
The lower minimum (upper maximum) bounds of the two inequalities are taken as the lower (upper) bounds of the entanglement fidelity.